# Untangling Lariats: Subgradient Following of Variationally Penalized Objectives

### v1

@@ 3 Derived Algorithms
In this section we apply the subgradient following framework to specific settings. We first show that the fused lasso is a simple instance. We also derive other variational (fused) penalties, including barrier penalties that were not analyzed before. We also show that the squared distance between ${\mathbf{x}}$ and $\mathbf{y}$ can be generalized at no additional computational cost to separable Bregman divergences.
@@ 3.1 Revisiting the Fused Lasso
To recover the Fused Lasso algorithm from the set following schema we choose, $h_{i}(x)=\frac{1}{2}(x-y_{i})^{2}$ and $g_{i}(\delta)=\lambda_{i}|\delta|$. Using basic calculus these choices imply that $\nabla h_{i}(x)=x-y_{i}$ and $g_{i}^{*}(\alpha)=\mathbf{1}_{|\alpha|\leq\lambda_{i}}$. For the domain of $g_{i}^{\star}$, namely $|\alpha|\leq\lambda_{i}$, its subgradient is,
$\partial g_{i}^{*}(\alpha)=\begin{cases}(-\infty,0]&\alpha=-\lambda_{i}\\ 0&|\alpha|<\lambda_{i}\\ [0,\infty)&\alpha=\lambda_{i}\end{cases}~{}.$
We next show that $\forall i\in[n]$, there exists a mapping $\mathsf{x}\mapsto\mathsf{a}_{i}(\mathsf{x})$, where $\mathsf{a}_{i}(\cdot)$ is a continuous non-decreasing piecewise linear function in $\mathsf{x}$, which serves as a succinct representation of $S_{i}=\{(\mathsf{x},\mathsf{a}_{i}(\mathsf{x}))\}$. The mapping takes the following recursive form,
$\mathsf{a}_{i}(\mathsf{x})=\big{[}\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}\big{]}_{-\lambda_{i}}^{+\lambda_{i}}~{}.$ (7)
First, let us define $a_{n}(\mathsf{x})\equiv 0$ which adheres with the definition of $S_{n}$ from previous section. Next, we show that the form of $S_{i}$ holds inductively by using the definitions of $S_{i}$ and $S_{i+\frac{1}{2}}$ from previous section. Assume that Eq. 7 holds true for $S_{i+1}$. The definition of the recursive map for $i+\frac{1}{2}$ implies that,
$S_{i+\frac{1}{2}}=\{(\mathsf{x},\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}):\mathsf{x}\in\mathsf{R}\}~{}.$
@@
we need to analyze $\Delta\mathsf{x}=\mathsf{x}-\mathsf{x}^{\prime}$ for the admissible set of values of $\mathsf{a}$ under the map $g_{i}^{*}$ which is $[-\lambda_{i},\lambda_{i}]$. Since $\mathsf{a}_{i+1}(\cdot)$ is continuous and non-decreasing, we have $\mathsf{a}_{i}^{\prime}(.)$ is strictly-increasing and bijective. Therefore, there exists two unique boundary values $z_{i}^{-}\;\text{s.t.}\;\mathsf{a}_{i}^{\prime}(z_{i}^{-})=-\lambda_{i}$ and $z_{i}^{+}\;\text{s.t.}\;\mathsf{a}_{i}^{\prime}(z_{i}^{+})=+\lambda_{i}$. For $\mathsf{a}=\mathsf{a}_{i}^{\prime}(\mathsf{x}^{\prime})$ such that $-\lambda_{i}<\mathsf{a}<\lambda_{i}$ we have $\partial g^{*}_{i}(\mathsf{a})=\{0\}$. Thus, $\mathsf{x}=\mathsf{x}^{\prime}$ and both are in $(z_{i}^{-},z_{i}^{+})$. We therefore get that the restriction of $S_{i}$ to $\mathsf{x}\in(z_{i}^{-},z_{i}^{+})$ yields,
@@
$=$ $\{(\mathsf{x},\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}):z_{i}^{-}<\mathsf{x}<z_{i}^{+}\}~{}.$
We last need to address the cases $\mathsf{a}=\pm\lambda_{i}$. For $\mathsf{a}=\mathsf{a}_{i}^{\prime}(z_{i}^{+})=\lambda_{i}$ we have $\partial g^{*}_{i}(\mathsf{a})=[0,\infty)$. Thus, for any $\mathsf{x}\geq z_{i}^{+}$ we have $\mathsf{x}-z_{i}^{+}\in\partial g^{*}_{i}(\lambda_{i})$ and therefore $(\mathsf{x},\lambda_{i})\in S_{i}$. Analogously, for $\mathsf{a}=-\lambda_{i}$ which implies that for $\mathsf{x}\leq z_{i}^{-}$ it holds that $\mathsf{x}-z_{i}^{-}\in\partial g^{*}_{i}(-\lambda_{i})\Rightarrow(\mathsf{x},-\lambda_{i})\in S_{i}$. We therefore get that the restriction of $S_{i}$ to $\mathsf{x}\geq z_{i}^{+}$ and $\mathsf{x}\leq z_{i}^{-}$ yields,
@@
$S_{i}\cap\{\mathsf{x}:\mathsf{x}\leq z_{i}^{-}\}$ $=$ $\{(\mathsf{x},-\lambda_{i}):\mathsf{x}\leq z_{i}^{-}\}~{}.$
@@
$S_{i}=\left\{\Big{(}\mathsf{x},\big{[}\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}\big{]}_{-\lambda_{i}}^{+\lambda_{i}}\Big{)}:\mathsf{x}\in\mathsf{R}\right\}~{}.$
@@ 3.2 Sparse Fused Lasso
$h_{i}(x)=\frac{1}{2}(x-y_{i})^{2}+\beta_{i}|x|~{}.$
and $g_{i}(\delta)=\lambda_{i}|\delta|$. This choice implies that
$\partial h_{i}(x)=\begin{cases}x-y_{i}+\beta_{i}&x>0\\ [-y_{i}-\beta_{i},-y_{i}+\beta_{i}]&x=0\\ x-y_{i}-\beta_{i}&x<0\end{cases}~{}.$
Generalizing the fused lasso, there exists a succinct representation for the sparse fused lasso of the form
$S_{i}=\{(\mathsf{x},\mathsf{a}_{i}^{+}(\mathsf{x})):\mathsf{x}>0\}\,\cup\,\{(0,\mathsf{a}):\mathsf{a}_{i}^{-}(0)\leq\mathsf{a}\leq\mathsf{a}_{i}^{+}(0)\}\,\cup\,\{(\mathsf{x},\mathsf{a}_{i}^{-}(\mathsf{x})):\mathsf{x}<0\}~{}.$
@@
$\mathsf{a}_{i}^{\pm}(\mathsf{x})=\big{[}\mathsf{a}_{i+1}^{\pm}(\mathsf{x})+\mathsf{x}-y_{i+1}\pm\beta_{i+1}\big{]}_{-\lambda_{i}}^{+\lambda_{i}}~{}.$
The sole value of that requires further attention is at $\mathsf{x}=0$. The admissible interval at zero is determined by the two boundary values, $\mathsf{a}_{i}^{-}(0)$ and $\mathsf{a}_{i}^{+}(0)$. To facilitate an efficient implementation, we store the value of $\Delta_{i}=\mathsf{a}_{i}^{+}(0)-\mathsf{a}_{i}^{-}(0)$. Before performing the rectification, we let $\Delta_{i}\leftarrow\Delta_{i-1}+2\beta_{i}$. Analogous to the fused lasso, we traverse the segments of $\mathsf{a}_{i}(\cdot)$ left-to-right until we reach the segment encapsulating $-\lambda_{i}$. The only difference is that we increase the value of $\mathsf{a}_{i}(\cdot)$ by $\Delta_{i}$ upon crossing the zero. Analogously, we decrease the value of $\mathsf{a}_{i}(\cdot)$ by $\Delta_{i}$ upon crossing the zero when traversing right-to-left. After obtaining the critical points $z_{i}^{\pm}$, we update $\Delta_{i}$ accordingly. In each iteration, we incur constant additional costs and thus the total runtime remains intact as $\mathcal{O}(n)$.
@@ 3.3 Bregman Divergences
Before proceeding with less explored fused regularizations, we present next a simple mean for incorporating convex losses for $h_{i}(\cdot)$ at zero additional computational cost. To remind the reader, given a strictly convex function $\phi:\mathbb{R}\to\mathbb{R}$, the Bregman divergence between $x\in\mathbb{R}$ and $y\in\mathbb{R}$ is defined as, $D_{\phi}(x,y)=\phi(x)-(\phi(y)+\nabla\phi(y)(x-y))$. Our generalization amount to setting, $h_{i}(x)=D_{\phi}(x,y_{i})$ and keeping $g_{i}(\delta)$ intact, thus retaining for the fused lasso $g_{i}(\delta)=\lambda_{i}|\delta|$. This choice for $h_{i}$ implies that $\nabla h_{i}(x)=\nabla\phi(x)-\nabla\phi(y_{i})$. Since $\phi(\cdot)$ is strictly convex $\nabla\phi(\cdot)$ is continuous and strictly increasing. Following the same line of derivation as above we get that the recursive mappings for each $\mathsf{a}_{i}(\cdot)$ takes the following form,
@@
This implies that $\mathsf{a}_{i}(.)$ is a piecewise linear as a function of $\nabla\phi(.)$. We thus follow the same procedure above whilst replacing $\mathsf{x}\mapsto\nabla\phi(x)$ for each linear segment in $\mathsf{a}_{i}$. This generalization to Bregman divergences is applicable throughout the remainder of the section. We keep using the squared error For brevity of our derivation in the sequel.
@@ 3.4 Isotonic Regression
$\partial g_{i}^{*}(\alpha)=\begin{cases}(-\infty,0]&\alpha=0\\ 0&\alpha>0\end{cases}~{}.$
@@
where $[z]_{+}=\max(z,0)$. The derivation follows the same line as above as there exists a unique boundary value $z_{i}^{-}\;\text{s.t.}\;\mathsf{a}_{i}^{\prime}(z_{i}^{-})=0$ and the resulting description of $S_{i}$ is obtained from the union of,
@@
$S_{i}\cap\{\mathsf{x}:\mathsf{x}\leq z_{i}^{-}\}$ $=$ $\{(\mathsf{x},0):\mathsf{x}\leq z_{i}^{-}\}~{}.$
@@
$g_{i}(\delta)=\bar{\lambda}_{i}[\delta]_{+}\,-\,\underline{\lambda}_{i}[-\delta]_{+}\,\mbox{ with }\,\bar{\lambda}_{i}\geq 0\,,\,\underline{\lambda}_{i}\leq 0~{}.$
@@
$\mathsf{a}_{i}(\mathsf{x})=\big{[}\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}\big{]}_{\underline{\lambda}_{i}}^{\bar{\lambda}_{i}}~{}.$
Setting $-\underline{\lambda}_{i}=\bar{\lambda}_{i}=\lambda_{i}$ yields the fused lasso whereas choosing $\underline{\lambda}_{i}=0$ and $\bar{\lambda}_{i}=\infty$ gives isotonic regression. In addition, choosing $\underline{\lambda}_{i}=0$ and $\bar{\lambda}_{i}=\lambda$ recovers the setting of [17]. To conclude this section we would like to note that isotonic regression can be used as the core for finding the closest vector to $\mathbf{y}$ subject to a $\leq k$-modality constraint. Due to submodularity of the objective the total runtime would be $\mathcal{O}(kn)$.
@@ 3.5 Fused Barriers
In certain data science applications the goal is to eliminate outlier points in an observed sequence rather than smooth or filter the sequence. We cast the problem as requiring the differences between consecutive points not to exceed prescribed maximal changes $b_{i}$ by establishing a series of barrier functions, $g_{i}(\delta)=\mathbf{1}_{|\delta|\leq\lambda_{i}}$. Each $b_{i}$ here can be thought of the maximum tolerance to changes and is not necessarily small and thus restrictive. For this choice of $g_{i}$ we get,
$\partial g_{i}^{*}(\alpha)=\begin{cases}-\lambda_{i}&\alpha<0\\ [-\lambda_{i},\lambda_{i}]&\alpha=0\\ \lambda_{i}&\alpha>0\end{cases}~{}.$ (9)
We next show that there exists a succinct representation of $S_{i}$ in the form of a piecewise linear map, $\mathsf{x}\mapsto\mathsf{a}_{i}(\mathsf{x})$ albeit being slightly more involved. First, let us reuse the intermediate function $\mathsf{a}_{i}^{\prime}$ from the derivation of the fused lasso algorithm, $\mathsf{a}_{i}^{\prime}(\mathsf{x})=\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}$. We denote by $\kappa_{\lambda}(\cdot)$ the Kvetsh operator (from Yiddish, literally “to squeeze, shrink”), colloquially called shrinkage,
$\kappa_{\lambda}(z)=\text{sign}(z)\,\big{[}|z|-\lambda\big{]}_{+}~{}.$
@@
Informally, after constructing the intermediate function $\mathsf{a}^{\prime}_{i}$ which is strictly monotone, we find the point $z_{i}$ where $\mathsf{a}^{\prime}_{i}$ crosses zero. We then insert an interval of zero slope whose center is $z_{i}$ and spans from $z_{i}-\lambda_{i}$ through $z_{i}+\lambda_{i}$.
We now provide the details of the derivation of $\mathsf{a}_{i}$. Since $z_{i}$ is the locus of the zero of $\mathsf{a}^{\prime}_{i}$ then for $\mathsf{x}^{\prime}>z_{i}$ we have $\mathsf{a}=\mathsf{a}_{i}^{\prime}(\mathsf{x}^{\prime})>0\Rightarrow\partial g_{i}^{*}(\mathsf{a})=\lambda_{i}$ based on Eq. 9. Thus, for any $\mathsf{x}=\mathsf{x}^{\prime}+\lambda_{i}>z_{i}+\lambda_{i}$ we have $(\mathsf{x},\mathsf{a}_{i}^{\prime}(\mathsf{x}-\lambda_{i}))\in S_{i}$. Analogously, for $\mathsf{x}<z_{i}-\lambda_{i}$ we have $(\mathsf{x},\mathsf{a}_{i}^{\prime}(\mathsf{x}+\lambda_{i}))\in S_{i}$. Last, for $(z_{i},\mathsf{a}^{\prime}_{i}(z_{i}))=(z_{i},0)$ we have $\partial g_{i}^{*}(0)=[-\lambda_{i},\lambda_{i}]$. Therefore, for any $\mathsf{x}$ such that $z_{i}-\lambda_{i}\leq\mathsf{x}\leq z_{i}+\lambda_{i}$ it holds that $\mathsf{x}-z_{i}\in\partial g_{i}^{*}(0)\Rightarrow(\mathsf{x},0)\in S_{i}$. The resulting $S_{i}$ is obtained from the union of,
@@
$S_{i}\cap\{\mathsf{x}\,:\,\mathsf{x}>z_{i}+\lambda_{i}\}$ $=$ $\{(\mathsf{x},\,\mathsf{a}_{i}^{\prime}(\mathsf{x}+\lambda_{i}))\,:\,\mathsf{x}>z_{i}+\lambda_{i}\}~{}.$
@@
Since each $\mathsf{a}_{i}(\cdot)$ is a piecewise linear function we can use the same algorithmic skeleton used for instance in the fused lasso. Alas, in order to construct $\mathsf{a}_{i}$ we need to traverse $\mathsf{a}^{\prime}_{i}$ and find the locus, $z_{i}$, of its zero. This search would require $\mathcal{O}(n)$ time and thus the total run time would amount to $\mathcal{O}(n^{2})$. We provide in Appendix A a describe more efficient implementation that require a more elaborated data structure in order to reduce the amortized runtime to $\mathcal{O}(n\log(n))$.

### v2

@@ 3 Derived Algorithms
In this section we apply the subgradient following framework to specific settings. We first show that the fused lasso is a simple instance. We next derive other variational (fused) penalties, including barrier penalties that were not analyzed before. We also show that the squared distance between ${\mathbf{x}}$ and $\mathbf{y}$ can be generalized at no additional computational cost to separable Bregman divergences.
@@ 3.1 Revisiting the Fused Lasso
To recover the Fused Lasso algorithm using the subgradient-following method we choose, $h_{i}(x)=\frac{1}{2}(x-y_{i})^{2}$ and $g_{i}(\delta)=\lambda_{i}|\delta|$. Using basic calculus these choices imply that $\nabla h_{i}(x)=x-y_{i}$ and $g_{i}^{*}(\alpha)=\mathbf{1}_{|\alpha|\leq\lambda_{i}}$. Over the domain of $g_{i}^{\star}$, namely $|\alpha|\leq\lambda_{i}$, its subgradient is,
$\partial g_{i}^{*}(\alpha)=\begin{cases}(-\infty,0]&\alpha=-\lambda_{i}\\ 0&|\alpha|<\lambda_{i}\\ [0,\infty)&\alpha=\lambda_{i}\end{cases}~{}~{}.$
We next show that $\forall i\in[n]$, there exists a mapping $\mathsf{x}\mapsto\mathsf{a}_{i}(\mathsf{x})$, where $\mathsf{a}_{i}(\cdot)$ is a continuous non-decreasing piecewise linear function in $\mathsf{x}$. This property serves as a succinct representation of $S_{i}=\{(\mathsf{x},\mathsf{a}_{i}(\mathsf{x}))\}$ which takes the following recursive form,
$\mathsf{a}_{i}(\mathsf{x})=\Big{[}\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}\Big{]}_{-\lambda_{i}}^{+\lambda_{i}}~{}~{}.$ (7)
To start, let us define $a_{n}(\mathsf{x})\equiv 0$ which adheres with the definition of $S_{n}$ from the previous section. We next show that the form of $S_{i}$ holds inductively, using the definitions of $S_{i}$ and $S_{i+\frac{1}{2}}$ from the previous section. Assume that Eq. (7) holds true for $S_{i+1}$. The definition of the recursive map for $i+\frac{1}{2}$ implies that,
$S_{i+\frac{1}{2}}=\{(\mathsf{x},\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}):\mathsf{x}\in\mathsf{R}\}~{}~{}.$
@@
we need to analyze $\Delta\mathsf{x}=\mathsf{x}-\mathsf{x}^{\prime}$ for the admissible set of values of $\mathsf{a}$ under the map $g_{i}^{*}$ which is $[-\lambda_{i},\lambda_{i}]$. Since $\mathsf{a}_{i+1}(\cdot)$ is continuous and non-decreasing, it implies that $\mathsf{a}_{i}^{\prime}(.)$ is strictly-increasing and bijective. Therefore, there exists two unique boundary values $z_{i}^{-}\;\text{s.t.}\;\mathsf{a}_{i}^{\prime}(z_{i}^{-})=-\lambda_{i}$ and $z_{i}^{+}\;\text{s.t.}\;\mathsf{a}_{i}^{\prime}(z_{i}^{+})=+\lambda_{i}$. For $\mathsf{a}=\mathsf{a}_{i}^{\prime}(\mathsf{x}^{\prime})$ such that $-\lambda_{i}<\mathsf{a}<\lambda_{i}$ we have $\partial g^{*}_{i}(\mathsf{a})=\{0\}$. Thus, $\mathsf{x}=\mathsf{x}^{\prime}$ and both are in $(z_{i}^{-},z_{i}^{+})$. We therefore get that the restriction of $S_{i}$ to $\mathsf{x}\in(z_{i}^{-},z_{i}^{+})$ yields,
@@
$=$ $\{(\mathsf{x},\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}):z_{i}^{-}<\mathsf{x}<z_{i}^{+}\}~{}~{}.$
We last need to address the cases $\mathsf{a}=\pm\lambda_{i}$. For $\mathsf{a}=\mathsf{a}_{i}^{\prime}(z_{i}^{+})=\lambda_{i}$ we have $\partial g^{*}_{i}(\mathsf{a})=[0,\infty)$. Thus, for any $\mathsf{x}\geq z_{i}^{+}$ we have $\mathsf{x}-z_{i}^{+}\in\partial g^{*}_{i}(\lambda_{i})$ and therefore $(\mathsf{x},\lambda_{i})\in S_{i}$. Analogously, for $\mathsf{a}=-\lambda_{i}$ which implies that for $\mathsf{x}\leq z_{i}^{-}$, it holds that $\mathsf{x}-z_{i}^{-}\in\partial g^{*}_{i}(-\lambda_{i})\Rightarrow(\mathsf{x},-\lambda_{i})\in S_{i}$. We therefore get that the restriction of $S_{i}$ to $\mathsf{x}\geq z_{i}^{+}$ and $\mathsf{x}\leq z_{i}^{-}$ yields,
@@
$S_{i}\cap\{\mathsf{x}:\mathsf{x}\leq z_{i}^{-}\}$ $=$ $\{(\mathsf{x},-\lambda_{i}):\mathsf{x}\leq z_{i}^{-}\}~{}~{}.$
@@
$S_{i}=\left\{\Big{(}\mathsf{x},\Big{[}\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}\Big{]}_{-\lambda_{i}}^{+\lambda_{i}}\Big{)}:\mathsf{x}\in\mathsf{R}\right\}~{}~{}.$
@@ 3.2 Sparse Fused Lasso
$h_{i}(x)=\frac{1}{2}(x-y_{i})^{2}+\beta_{i}|x|~{}\mbox{ and }~{}g_{i}(\delta)=\lambda_{i}|\delta|~{}~{}.$
These choices imply that
$\partial h_{i}(x)=\begin{cases}x-y_{i}+\beta_{i}&x>0\\ [-y_{i}-\beta_{i},-y_{i}+\beta_{i}]&x=0\\ x-y_{i}-\beta_{i}&x<0\end{cases}~{}~{}.$
Generalizing the fused lasso, the succinct representation for the sparse fused lasso is of the form,
$S_{i}=\{(\mathsf{x},\mathsf{a}_{i}^{+}(\mathsf{x})):\mathsf{x}>0\}\,\cup\,\{(0,\mathsf{a}):\mathsf{a}_{i}^{-}(0)\leq\mathsf{a}\leq\mathsf{a}_{i}^{+}(0)\}\,\cup\,\{(\mathsf{x},\mathsf{a}_{i}^{-}(\mathsf{x})):\mathsf{x}<0\}~{}~{}.$
@@
$\mathsf{a}_{i}^{\pm}(\mathsf{x})=\Big{[}\mathsf{a}_{i+1}^{\pm}(\mathsf{x})+\mathsf{x}-y_{i+1}\pm\beta_{i+1}\Big{]}_{-\lambda_{i}}^{+\lambda_{i}}~{}~{}.$
The sole value that requires further attention is at $\mathsf{x}=0$. The admissible interval at zero is determined by the two boundary values, $\mathsf{a}_{i}^{-}(0)$ and $\mathsf{a}_{i}^{+}(0)$. To facilitate an efficient implementation, we store the value of $\Delta_{i}=\mathsf{a}_{i}^{+}(0)-\mathsf{a}_{i}^{-}(0)$. Prior to rectification, we update $\Delta_{i}\leftarrow\Delta_{i-1}+2\beta_{i}$. Analogous to the fused lasso, we traverse the segments of $\mathsf{a}_{i}(\cdot)$ left-to-right until we reach the segment encapsulating $-\lambda_{i}$. The only difference is that we increase the value of $\mathsf{a}_{i}(\cdot)$ by $\Delta_{i}$ upon crossing the zero. Analogously, we decrease the value of $\mathsf{a}_{i}(\cdot)$ by $\Delta_{i}$ upon crossing the zero when traversing right-to-left. After obtaining the critical points $z_{i}^{\pm}$, we update $\Delta_{i}$ accordingly. In each iteration, we incur constant additional costs and thus the total runtime remains intact as $\mathcal{O}(n)$.
@@ 3.3 Bregman Divergences
Before proceeding with nascent total variation regularizers, we now describe a simple way for incorporating convex losses for $h_{i}(\cdot)$ at zero additional computational cost. To remind the reader, given a strictly convex function $\phi:\mathbb{R}\to\mathbb{R}$, the Bregman divergence between $x\in\mathbb{R}$ and $y\in\mathbb{R}$ is defined as, $D_{\phi}(x,y)=\phi(x)-(\phi(y)+\nabla\phi(y)(x-y))$. Our generalization amount to setting, $h_{i}(x)=D_{\phi}(x,y_{i})$ and keeping $g_{i}(\delta)$ intact, thus retaining for the fused lasso $g_{i}(\delta)=\lambda_{i}|\delta|$. This choice for $h_{i}$ implies that $\nabla h_{i}(x)=\nabla\phi(x)-\nabla\phi(y_{i})$. Since $\phi(\cdot)$ is strictly convex, $\nabla\phi(\cdot)$ is continuous and strictly increasing. Following the same line of derivation as above we get that the recursive mappings for each $\mathsf{a}_{i}(\cdot)$ takes the following form,
@@
This implies that $\mathsf{a}_{i}(.)$ is a piecewise linear function of $\nabla\phi(.)$. We thus follow the same procedure above whilst replacing $\mathsf{x}\mapsto\nabla\phi(x)$ for each linear segment in $\mathsf{a}_{i}$. This generalization using a Bregman divergence is applicable throughout the remainder of the section. We keep using the squared error for brevity of our derivations.
@@ 3.4 Isotonic Regression
$\partial g_{i}^{*}(\alpha)=\begin{cases}(-\infty,0]&\alpha=0\\ 0&\alpha>0\end{cases}~{}~{}.$
@@
where $[z]_{+}=\max(z,0)$. The derivation follows the same lines as above as there exists a unique boundary value $z_{i}^{-}\;\text{s.t.}\;\mathsf{a}_{i}^{\prime}(z_{i}^{-})=0$. The resulting description of $S_{i}$ is obtained from the union of,
@@
$S_{i}\cap\{\mathsf{x}:\mathsf{x}\leq z_{i}^{-}\}$ $=$ $\{(\mathsf{x},0):\mathsf{x}\leq z_{i}^{-}\}~{}~{}.$
@@
$g_{i}(\delta)=\bar{\lambda}_{i}[\delta]_{+}\,-\,\underline{\lambda}_{i}[-\delta]_{+}\,\mbox{ with }\,\bar{\lambda}_{i}\geq 0\,,\,\underline{\lambda}_{i}\leq 0~{}~{}.$
@@
$\mathsf{a}_{i}(\mathsf{x})=\big{[}\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}\big{]}_{\underline{\lambda}_{i}}^{\bar{\lambda}_{i}}~{}~{}.$
Setting $-\underline{\lambda}_{i}=\bar{\lambda}_{i}=\lambda_{i}$ yields the fused lasso whereas choosing $\underline{\lambda}_{i}=0$ and $\bar{\lambda}_{i}=\infty$ provides a template for isotonic regression. In addition, choosing $\underline{\lambda}_{i}=0$ and $\bar{\lambda}_{i}=\lambda$ recovers the setting of [28]. To conclude this section we would like to note that isotonic regression can be used as the core for finding the closest vector to $\mathbf{y}$ subject to a $\leq k$-modality constraint. Due to submodularity of the objective the total runtime would be $\mathcal{O}(kn)$.
@@ 3.5 Fused Barriers
For certain data analysis applications the goal is to eliminate outlier points in an observed sequence rather than smooth or filter the sequence. We cast the problem as requiring the differences between consecutive points not to exceed prescribed maximal changes $b_{i}$ by establishing a series of barrier functions, $g_{i}(\delta)=\mathbf{1}_{|\delta|\leq\lambda_{i}}$. Each $b_{i}$ here can be thought of as the maximum tolerance to changes and is not necessarily small and thus restrictive. For this choice of $g_{i}$ we get,
$\partial g_{i}^{*}(\alpha)=\begin{cases}-\lambda_{i}&\alpha<0\\ [-\lambda_{i},\lambda_{i}]&\alpha=0\\ \lambda_{i}&\alpha>0\end{cases}~{}~{}.$ (9)
Albeit being slightly more involved, we next show that there exists a succinct representation of $S_{i}$ in the form of a piecewise linear map, $\mathsf{x}\mapsto\mathsf{a}_{i}(\mathsf{x})$. First, let us reuse the intermediate function $\mathsf{a}_{i}^{\prime}$ from the derivation of the fused lasso algorithm, $\mathsf{a}_{i}^{\prime}(\mathsf{x})=\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}$. We denote by $\kappa_{\lambda}(\cdot)$ the Kvetsh operator (from Yiddish, literally “to squeeze, shrink”), colloquially called shrinkage,
$\kappa_{\lambda}(z)=\text{sign}(z)\,\big{[}|z|-\lambda\big{]}_{+}~{}~{}.$
@@
Informally, after constructing the intermediate function $\mathsf{a}^{\prime}_{i}$, which is strictly monotone, we find the point $z_{i}$ where $\mathsf{a}^{\prime}_{i}$ crosses zero. We then insert an interval of zero slope whose center is $z_{i}$ and its support spans from $z_{i}-\lambda_{i}$ through $z_{i}+\lambda_{i}$.
We now provide the details of the derivation of $\mathsf{a}_{i}$. Since $z_{i}$ is the locus of the zero of $\mathsf{a}^{\prime}_{i}$ then for $\mathsf{x}^{\prime}>z_{i}$ we have $\mathsf{a}=\mathsf{a}_{i}^{\prime}(\mathsf{x}^{\prime})>0\Rightarrow\partial g_{i}^{*}(\mathsf{a})=\lambda_{i}$ based on Eq. (9). Thus, for any $\mathsf{x}=\mathsf{x}^{\prime}+\lambda_{i}>z_{i}+\lambda_{i}$ we have $(\mathsf{x},\mathsf{a}_{i}^{\prime}(\mathsf{x}-\lambda_{i}))\in S_{i}$. Analogously, for $\mathsf{x}<z_{i}-\lambda_{i}$ we have $(\mathsf{x},\mathsf{a}_{i}^{\prime}(\mathsf{x}+\lambda_{i}))\in S_{i}$. Last, for $(z_{i},\mathsf{a}^{\prime}_{i}(z_{i}))=(z_{i},0)$ we have $\partial g_{i}^{*}(0)=[-\lambda_{i},\lambda_{i}]$. Therefore, for any $\mathsf{x}$ such that $z_{i}-\lambda_{i}\leq\mathsf{x}\leq z_{i}+\lambda_{i}$ it holds that $\mathsf{x}-z_{i}\in\partial g_{i}^{*}(0)\Rightarrow(\mathsf{x},0)\in S_{i}$. The resulting $S_{i}$ is obtained from the union of,
@@
$S_{i}\cap\{\mathsf{x}\,:\,\mathsf{x}>z_{i}+\lambda_{i}\}$ $=$ $\{(\mathsf{x},\,\mathsf{a}_{i}^{\prime}(\mathsf{x}+\lambda_{i}))\,:\,\mathsf{x}>z_{i}+\lambda_{i}\}~{}~{}.$
@@
Since each $\mathsf{a}_{i}(\cdot)$ is a piecewise linear function we can employ the same algorithmic skeleton used for the fused lasso. Alas, in order to construct $\mathsf{a}_{i}$ we need to traverse $\mathsf{a}^{\prime}_{i}$ and find the locus, $z_{i}$, of its zero. This search would require $\mathcal{O}(n)$ time and thus the total run time would amount to $\mathcal{O}(n^{2})$. We provide in Appendix A a description of a more efficient implementation which uses a more elaborated data structure in order to reduce the amortized runtime to $\mathcal{O}(n\log(n))$.
@@
We conclude the section with a short discussion and illustration of the characteristics of the variational penalties of the fused lasso by contrasting it with barrier constraints on the variation. To this end, we generated a sequence based on a quadratic function, $y_{i}=(i-50)^{2}/100$. We then contaminated the sequence with noise. We first added i.i.d noise to each element sampled from the uniform distribution over $[-\epsilon,\epsilon]$. We next added “shock” noise by replacing the value of a few elements chosen at random with a small negative number. The generating and noisy sequences are plotted on the left hand-side of Figure 5. We then solved least squares approximation for a range of variational penalties and constraints by varying $\lambda$. For each solution ${\mathbf{x}}(\lambda)$ we calculated the $\ell_{1}$ and $\ell_{\infty}$ variation norm. The results are shown on the right hand-side of Figure 5. The dashed two-sided arrows designate solutions attaining the same approximation error by the two methods. It is evident that for the same error the fused lasso naturally obtains a lower $\ell_{1}$ variation norm and it is more resilient to the uniform noise. In contrast, posing a barrier constraint on the variation provides a uniformly lower $\ell_{\infty}$ norm of the variation and is more resilient to shock noise.
Figure 5: Comparison least-squares sequence approximation with variational penalty of the fused lasso (FL) versus barrier constraints (FB).